Show that equivalence modulo n in Z+ is an equivalence relation.


1
Expert's answer
2021-02-01T19:03:58-0500

Solution:

To show: Congruence modulo in Z is an equivalence relation.

Reflexive:

Given, equal numbers are congruent.

"\forall x, y, z \in \mathbb{Z}: x=y \Longrightarrow x \equiv y(\bmod z)"

so it implies that:

"\forall x \in \mathbb{Z}: x \equiv x(\bmod z)"

and so congruence modulo z is reflexive.

Symmetric:

"\begin{aligned}
x & \equiv y \bmod z \\
\Rightarrow \quad x-y &=k z \\
\Rightarrow \quad y-x &=(-k) z \\
\Rightarrow y & \equiv x \bmod z
\end{aligned}"

So congruence modulo z is symmetric.

Transitive:

"\begin{aligned}
x_{1} & \equiv x_{2} \quad(\bmod z) \\
x_{2} & \equiv x_{3} \quad(\bmod z) \\
\left(x_{1}-x_{2}\right) &=k_{1} z
\end{aligned}"

"\begin{aligned}
\wedge &\left(x_{2}-x_{3}\right)=k_{2} z \\
\Rightarrow &\left(x_{1}-x_{2}\right)+\left(x_{2}-x_{3}\right)=\left(k_{1}+k_{2}\right) z \\
\Rightarrow &\left(x_{1}-x_{3}\right)=\left(k_{1}+k_{2}\right) z
\end{aligned}
x_{1} \equiv x_{3} \quad(\bmod z)"

"\Rightarrow x_{1} \equiv x_{3} \quad(\bmod z)"

So congruence modulo z is transitive.

Thus, it is an equivalence relation.

Hence, proved.